### v1

# AdaCoT: Rethinking Cross-Lingual Factual Reasoning through Adaptive Chain-of-Thought

Abstract
Large language models (LLMs) have shown impressive multilingual capabilities through pretraining on diverse corpora. While these models show strong reasoning abilities, their performance varies significantly across languages due to uneven training data distribution. Existing approaches using machine translation, and extensive multilingual pretraining and cross-lingual tuning face scalability challenges and often fail to capture nuanced reasoning processes across languages. In this paper, we introduce AdaCoT (Adaptive Chain-of-Thought), a framework that enhances multilingual reasoning by dynamically routing thought processes through intermediary “thinking languages” before generating target-language responses. AdaCoT leverages a language-agnostic core and incorporates an adaptive, reward-based mechanism for selecting optimal reasoning pathways without requiring additional pretraining. Our comprehensive evaluation across multiple benchmarks demonstrates substantial improvements in both factual reasoning quality and cross-lingual consistency, with particularly strong performance gains in low-resource language settings. The results suggest that adaptive reasoning paths can effectively bridge the performance gap between high and low-resource languages while maintaining cultural and linguistic nuances.
{CJK}
UTF8gbsn
AdaCoT: Rethinking Cross-Lingual Factual Reasoning through
Adaptive Chain-of-Thought
Xin Huang, Tarun Kumar Vangani, Zhengyuan Liu, Bowei Zou,  Ai Ti Aw Institute for Infocomm Research (I2R), A*STAR, Singapore {huangx2, vangani_tarun_kumar, liuz2, zou_bowei, aaiti}@i2r.a-star.edu.sg
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable reasoning capabilities Achiam et al. (2023), yet their performance varies significantly across languages, with a notable bias toward high-resource languages like English. Moreover, this limits the effectiveness and accessibility of LLMs for speakers of other languages, despite increasing worldwide demand for AI systems that can serve diverse linguistic communities. While machine translation offers a straightforward solution, it proves inadequate for developing robust multilingual reasoning capabilities, often introducing artifacts and failing to capture nuanced logical processes across linguistic boundaries.
Previous work has focused primarily on data-level improvements through training on large-scale, balanced multilingual corpora, including continual pretraining Xu et al. (2024); Yang et al. (2023), multilingual instruction fine-tuning Li et al. (2023); Zhang et al. (2024), and cross-lingual instruction fine-tuning on translation pairs Zhang et al. (2023); Zhu et al. (2023). Recent efforts have explored representation-level solutions by aligning multilingual representations Li et al. (2024a, b), building on evidence that consistent multilingual representations enhance knowledge transfer Tang et al. (2022). However, these methods typically require extensive training data or fail to deliver uniform improvements across languages Zhu et al. (2023); Li et al. (2024a).
Studies on language representation subspaces in LLMs demonstrate that LLMs have a language-agnostic reasoning core that can be exploited to improve multilingual performance without extensive cross-lingual pretraining Tang et al. (2024); Zhao et al. (2024). However, while this core enables basic reasoning capabilities, the uneven distribution of cultural and domain-specific knowledge across languages leads to significant performance gaps in factual reasoning tasks. This challenge is particularly acute for low-resource languages, which receive limited exposure during training, and a framework leveraging language-specific strengths while preserving cultural context is needed. Building on this insight, we introduce AdaCoT, a novel framework for optimizing multilingual reasoning in LLMs through adaptive chain-of-thought. AdaCoT strategically routes reasoning steps through intermediate “thinking languages” before generating target-language outputs. Our key insight is that certain languages may be more effective for specific categories of reasoning tasks, whether due to structural characteristics, cultural context, or the model’s training distribution. For instance, languages with rich logical connectives may perform better in deductive reasoning, whereas languages with extensive mathematical vocabulary could be more adept at quantitative tasks.
The AdaCoT framework operates on two key principles: (1) Dynamic Routing Optimization, which learns to select the most effective intermediate languages based on task characteristics and performance history; and (2) Cross-Lingual Knowledge Integration, which synthesizes insights from multiple linguistic perspectives to produce more robust final outputs.
Our approach incorporates an adaptive reward-based mechanism that dynamically selects optimal reasoning pathways, learning to route queries through appropriate intermediate languages based on performance feedback. This selective approach allows the model to leverage the strengths of different languages while maintaining computational efficiency, as it requires no additional pretraining on large multilingual corpora. The reward mechanism considers factors such as reasoning consistency, answer accuracy, and linguistic fluency to continuously refine its routing decisions.
Empirical evaluation across multiple benchmark datasets, including multilingual Truthful-QA, CrossAlpaca-Eval 2.0, Cross-MMLU and Cross-LogiQA, demonstrates significant improvements in reasoning quality and consistency across diverse languages. Our analysis reveals particularly strong gains in low-resource languages when reasoning is routed through linguistically related high-resource languages, suggesting the existence of transferable reasoning patterns within language families. Furthermore, we observe that the adaptive routing mechanism learns meaningful patterns, such as preferring languages with precise technical vocabularies for scientific reasoning tasks and those with rich emotional expression for sentiment-based tasks. These findings not only validate the effectiveness of AdaCoT but also provide insights into the nature of multilingual reasoning in LLMs. The success of language-specific routing suggests that different languages may encode different cognitive patterns, and leveraging these differences can lead to more robust and nuanced reasoning capabilities. Our work opens new avenues for research into language-aware neural architectures and cross-lingual knowledge transfer mechanisms.
2 Related Work
Large language models (LLMs) exhibit multilingual capabilities through pretraining on diverse corpora Bang et al. (2023). While primarily developed for resource-rich languages such as English, French, and Chinese, LLMs have shown unexpected proficiency across languages Dubey et al. (2024). However, uneven training data distribution at language level leads to performance disparities between high and low-resource languages Qi et al. (2023); Wang et al. (2024a); Chua et al. (2024).
Effective multilingual alignment requires careful design of both pretraining and instruction tuning strategies Gao et al. (2024). Recent approaches include multilingual contrastive learning with cross-lingual instruction tuning Li et al. (2024a) and word-level alignment before code-switched pretraining Li et al. (2024b). Translation parallel pairs have proven particularly effective Zhang et al. (2023); Zhu et al. (2024); Mu et al. (2024); Lin et al. (2024). Notable advances include two-stage training combining extensive monolingual data with high-quality translation datasets Xu et al. (2024), integration of multilingual translation and task instructions Zhu et al. (2023), and self-distillation methods transferring capabilities from high to low-resource languages Zhang et al. (2024). While Cross-lingual Expert Language Models (X-ELM) train independently on multilingual corpus subsets Blevins et al. (2024), native language prompting remains crucial for culture-specific tasks requiring nuanced understanding Liu et al. (2024).
For cross-lingual reasoning, multilingual instruction tuning with English as the default thinking language improves reasoning consistency on math problems Lai and Nissim (2024). Moreover, layer transplantation between language and math experts, fine-tuned from a common pretrained model, enhances mathematical performance in target languages Bandarkar et al. (2024).
Recent studies on language processing dynamics reveal that English-centric LLMs use English as an internal pivot language Wendler et al. (2024), while non-English-centric models leverage their primary training language Zhong et al. (2024). Analysis of language-agnostic knowledge neurons and cross-lingual neural patterns suggests that aligning multilingual representations enhances knowledge transfer Wang et al. (2024c); Tang et al. (2024); Cao et al. (2024).
3 Methodology
We propose a novel framework, AdaCoT, for optimizing multilingual reasoning in LLMs through adaptive chain-of-thought. The core concept of AdaCoT is to leverage intermediate reasoning steps in one or more auxiliary “thinking languages” to improve the final output quality in a target language. This approach builds on the observation that, although LLM reasoning is largely language-agnostic, model performance can vary significantly across different languages due to disparities in training data, linguistic features, and alignment.
AdaCoT leverages the dynamic integration of reasoning tokens in thinking languages. The thinking languages are selected based on their linguistic similarity, resource availability, or alignment with the target language. The reasoning tokens in auxiliary languages guide the model toward generating more accurate and contextually appropriate responses in the target language. For instance, dominant or well-aligned languages with extensive resources (e.g., English) may serve as mediators for reasoning tasks in underrepresented languages. To this end, we propose a dual-pathway mechanism, where Cross-Lingual Chain-of-Thought leverages chain-of-thought reasoning in auxiliary languages to enhance output quality, while Direct Generation bypasses intermediate languages for efficient generation in well-supported target languages or contexts requiring minimal cross-linguistic reasoning.
The pathway selection is optimised through a reward-based fine-tuning process, making the model dynamically select the most effective reasoning strategy based on the input context. The optimisation process is to train the base model with reward signals derived from a GPT-based reward LLM, which the model learns to balance the trade-offs between intermediate and direct reasoning strategies. It leads to improved performance across diverse languages and contexts. An overview of AdaCoT is illustrated in Figure 1 and 2.
Figure 1: Training Data Generation for AdaCoT.
Figure 2: Inference without and with AdaCoT.
3.1 Candidate Response Generation
During the training stage, given a user query in a specific language, AdaCoT first leverages an LLM to translate the input into multiple primary languages, including the original input language and three auxiliary languages: English, Chinese, and Indonesian. Such multilingual translation step forces the model to reason with diverse linguistic contexts. Based on this setup, we employ two distinct reasoning strategies to generate a diverse pool of candidate outputs:
  • Cross-Lingual Chain-of-Thought. This approach uses intermediate reasoning steps in auxiliary languages to guide the creation of the final response. We begin by translating the original prompt into auxiliary languages such as English, Chinese, and Indonesian, then instruct the LLM to generate responses for each translated version. Because the internal knowledge of the model can vary between languages, these responses may differ and reflect unique linguistic and cultural nuances. To synthesize these diverse outputs into a coherent final response that both follows the original instructions of the prompt and preserves the underlying knowledge of each intermediate response, we employ a teacher model (e.g., GPT-4o) in each of the primary languages. Formally, given a prompt $P_{l}$ in language $l$ and an intermediate response $I_{t}$ in auxiliary language $t$, we instruct the teacher model to produce a final response $R_{l}^{t}$ in language $l$. The aim is for $R_{l}^{t}$ to keep the semantic meaning of $I_{t}$ while adhering to the original instruction $P_{l}$.
  • Direct Generation. Direct Generation creates responses in the intended language without relying on any auxiliary languages. Given a prompt $P^{l}$ in language $l$, the model directly produces the response $R_{l}^{l}$ in the same language. This strategy is especially useful if the model is already strong in that language or in situations where using multiple languages might degrade the performance, such as linguistic dependent tasks like writing poetry.
The algorithm 1 depicts the training pair generation process. With these two strategies, AdaCoT generates diverse responses across multiple language pathways, which can then be ranked by a powerful LLM. Because certain knowledge may not be available or shared in all languages, the approach helps reduce the risk of factual hallucinations.
Algorithm 1 Candidate Response Generation for AdaCoT
0:  Input prompt $P_{l}$ in the target language $l$; set of auxiliary languages $\theta$; a Base LLM for generation; a Teacher Model for response synthesis.
0:  A set of candidate responses $\{\,R_{l}^{t}\mid t\in\theta\}$, each reflecting reasoning in a primary language $t$.
1:  Step 1: Multilingual Prompt Translation
2:  For each auxiliary language $t\in\theta$:
3:   Translate $P_{l}$ into $t$, yielding $P_{l}^{t}$. If $t=l$, simply reuse $P_{l}.$
4:  Step 2: Cross-Lingual CoT & Direct Generation
5:  for each language $t$ in $\theta$ do
6:     if $t\neq l$ then
7:        Cross-Lingual CoT:
8:         1. Generate an intermediate response $I_{t}$ reasoning in language $t$ from the prompt $P_{l}^{t}$ using the Base LLM.
9:         2. Instruct the Teacher Model to produce $R_{l}^{t}$ in the target language $l$ using both the original prompt $P_{l}$ and the intermediate response $I_{t}$ while maintaining the semantic meaning of $I_{a}$ and adhere to the original instruction $P_{l}$.
10:     else
11:        Direct Generation:
12:         Use the Base LLM to generate $R_{l}^{l}$ directly from $P_{l}$.
13:     end if
14:  end for
3.2 Candidate Response Ranking
We score the diverse responses in different language pathways using an LLM as a judge, then select the optimal auxiliary language or opt for direct generation based on the highest score. Given an input prompt $P_{l}$ in language $l$ and a set of possible language pathways $\theta$, the LLM produces an output $R_{l}^{t}$ of corresponding language $l$ by thinking in auxiliary language $t\in\theta$. The reward LLM then provides a text-based score to rank responses generated in different languages or directly without intermediate reasoning.
For this work, we use GPT-4o as the reward and teacher LLM. We design a special prompt $p_{s}$ outlining four metrics that assess different dimensions of response quality, and feed each metric into the reward LLM. The final score is generated from the reward LLM considering all evaluation criteria including factual incorrectness, response hallucinations, repetition, and adherence to the instructions. We use a Likert-like scale ranging from 0 to 10 as the final ranking score.
$S_{t}=GPT([p_{s},P_{l},R_{l}^{t}])$
3.3 AdaCoT Fine-Tuning
We use the judging scores $S_{t}$ to guide the selection of optimal reasoning actions, maximizing the likelihood of choosing the pathway with the highest score across various primary thinking languages.
We fine-tune only on instances where $S_{t}>=9$ allowing the model to learn from reasoning strategies that yield the highest quality outputs. For each training example, we select the optimal reasoning pathway that attains the highest score, whether it involves reasoning through an intermediate language or directly generating in the target language. Note that both the intermediate response $I_{t}$ and the final response $R_{l}^{t}$ are generated by the LLM itself rather than using labels from external dataset, to prevent possible potential knowledge forgetting during fine-tuning. The fine-tuning process allows the model to dynamically adapt different thinking process based on contextual requirements of input prompts.
Training examples are formatted with special prompts to represent the reasoning pathways. For a input query $P_{l}$ with a corresponding final answer $R_{l}^{t}$, if intermediate reasoning $I_{t}$ is included, the training example is structured as:
$P_{l}\ [P_{l}^{q}]\ I_{t}\ [P_{l}^{r}]\ R_{l}^{t}$
We use $P_{l}^{q}$ to denote a special prompt that separates intermediate reasoning tokens from the query, and $P_{l}^{r}$ to separate the final answer from the intermediate tokens. Note that $R_{l}^{t}$ is not a direct translation of the intermediate reasoning tokens. Rather, $R_{l}^{t}$ is generated using $I_{t}$ as an in-context learning example. The distinction allows the model to generate responses that adhere to the user query, rather than generating responses with fixed language that could be contextually inappropriate if the response requires multiple languages.
When it’s direct generation of the answer without intermediate thinking, the format simplifies to:
$P_{l}\ [P_{l}^{q}]\ R_{l}^{l}$
We also set attention mask of $P_{l}$ to 0 during the fine-tuning process to force the model only predicts correct reasoning pathways and final responses given an input prompt.
4 Experimentation
4.1 Experiment Setup
4.1.1 Base Models
The AdaCoT is applicable to various open-source models. For this experiment, we use LLaMA3.1-8B-Instruct Dubey et al. (2024) and Qwen2.5-7B-Instruct Qwen et al. (2025) as the base models. LLaMA3.1-8B is particularly strong in English reasoning tasks and Qwen2.5-7B excels at reasoning in both English and Chinese. Thus we can explore AdaCoT’s impact in base LLMs that possess high-resource knowledge in multiple languages.
4.1.2 Primary Thinking Languages
We select English, Chinese, and Indonesian as intermediate reasoning languages. We choose English and Chinese for their extensive knowledge bases and linguistic resources, as well as the often unshared or misaligned knowledge between them, which can create performance gaps when transferring knowledge. We also include Indonesian as a low-resource language to evaluate AdaCoT’s ability to transfer knowledge from high-resource languages and reduce performance disparities.
4.1.3 Training Datasets
To determine best reasoning pathway for the base LLM, we need to explore diverse instructions and cover as many input instruction types as possible. To this end, we use 1 million English OpenHermes 2.5 Teknium (2023) and 1.1 million chinese instruction dataset from Firefly Yang (2023) as our base dataset. We further augment the dataset by leveraging GPT-4o model to translate the datasets which covers all primary thinking languages including English, Chinese, and Indonesian. We use only the instructions of these datasets, and leverage algorithm 1 and the Candidate Response Ranking method to generate the final dataset for fine-tuning.
4.2 Evaluation Datasets
4.2.1 Multilingual TruthfulQA
Multilingual TruthfulQA Lai et al. (2023) is designed to assess the truthfulness of LLMs in multilingual settings. It includes parallel questions translated from original TruthfulQA Lin et al. (2021) into 31 diverse languages using GPT-3.5-turbo.
4.2.2 CrossAlpaca-Eval 2.0
The CrossAlpaca-Eval 2.0 dataset is an open-ended question-answering dataset with paralleled instruction pairs in English, Chinese, and Indonesian languages from Alpaca-Eval 2.0 Dubois et al. (2024) by translating the original English instruction pairs into these languages via GPT-4o. We choose this dataset in order to understand the effectiveness of AdaCoT in diverse tasks.
4.2.3 Cross-MMLU and Cross-LogiQA
Model ar bn ca da de en es eu fr gu hi
LLaMA3.1-8B 51.05 40.79 52.76 48.82 53.42 62.06 54.74 38.42 50.92 38.55 43.42
LLaMA3.1-8B-AdaCoT 57.24 48.68 56.97 57.89 60.92 64.50 61.71 53.42 58.55 47.89 51.45
Qwen2.5-7B 55.66 43.29 54.34 54.21 56.71 61.4 58.29 32.37 52.63 37.24 46.45
Qwen2.5-7B-AdaCoT 58.16 38.29 55.79 55.39 59.21 62.8 59.74 48.42 58.68 34.74 45.00
Model hr hu hy id it kn ml mr ne nl pt
LLaMA3.1-8B 50.26 48.29 36.18 53.42 54.08 38.55 38.95 53.16 46.32 52.89 57.63
LLaMA3.1-8B-AdaCoT 56.05 52.76 49.34 60.66 62.11 43.03 42.89 47.37 52.50 61.58 59.08
Qwen2.5-7B 50.79 46.32 28.68 58.29 59.47 26.45 25.00 31.71 36.84 58.16 63.03
Qwen2.5-7B-AdaCoT 50.53 46.97 27.24 58.68 60.92 26.97 28.29 33.95 37.76 59.47 63.03
Model ro ru sk sr sv ta te uk vi zh
LLaMA3.1-8B 52.50 50.79 45.79 47.63 52.76 45.53 35.53 50.53 52.50 55.00
LLaMA3.1-8B-AdaCoT 58.29 60.26 53.03 56.58 59.74 45.39 42.63 60.13 58.29 60.79
Qwen2.5-7B 57.37 59.61 48.55 52.63 55.53 24.87 26.84 52.89 57.63 61.32
Qwen2.5-7B-AdaCoT 56.97 60.66 54.08 52.24 55.26 30.00 25.00 60.53 56.05 62.41
Table 1: Multilingual performance comparison on the mTruthfulQA dataset.
For logical reasoning evaluation, we use Cross-MMLU and Cross-LogiQA Wang et al. (2024b).
Cross-MMLU: Cross-MMLU is derived from the MMLU Hendrycks et al. (2021) dataset. The dataset contains parallel questions from MMLU in English, Chinese and Indonesian, and each language comprising 150 questions. The original English questions were initially translated using Google Translate, and have been edited and filtered by native speakers to ensure accurate translations and guarantee semantically equitable evaluations across all languages.
Cross-LogiQA: Cross-LogiQA is adapted from the LogiQA2.0 Liu et al. (2023) dataset, a benchmark for evaluating logical reasoning abilities in natural language understanding. The dataset consists of 176 questions per language across English, Chinese and Indonesian languages, developed using the same translation and verification process as Cross-MMLU to produce parallel versions in all languages. This dataset allows us to examine whether models can perform consistent logical deductions across different linguistic representations of the same logical problem.
4.2.4 Evaluation Metrics
Our evaluation consists of two primary evaluation metrics. For open-ended question-answering tasks in the CrossAlpaca-Eval 2.0 dataset, we use LLM-as-a-judge for the comprehensive evaluation of model responses. Recent work such as Zheng et al. (2024) finds that reward LLMs like GPT-4 align well with human preferences which archives over 80% agreement, therefore providing scalable and consistent evaluations of model outputs. We employ GPT-4o as judging model, we use likert scale from 0 to 10 to score model outputs based on criteria including correctness, coherence, and adherence to instructions.
For multiple-choice question tasks in the multilingual TruthfulQA, Cross-MMLU and Cross-LogiQA datasets, we implement an answer selection evaluation method. In this method, the evaluation LLM is prompted to select the correct answer from a set of labeled options provided in the model’s output. The model’s selections are then compared to the ground-truth answers to calculate accuracy. We use instructed version of gemma-2-27b Team et al. (2024) as the evaluator. This method provides an accurate evaluation of the model’s response in different styles and better handles cases where the model produces chain-of-thought tokens. Note that for AdaCoT models, we remove the intermediate thinking tokens before evaluation to ensure fair evaluations.
We additionally evaluate cross-lingual consistency $C$ for Cross-MMLU and Cross-LogiQA to measure the consistency of the answers regardless of whether answer is correct or not. Let $Q$ be the total number of questions, $L$ be the number of languages, and $\hat{y}_{l,q}$ be the model’s prediction for language $l$ on question $q$. We measure the consistency for all 3 languages using following formula:
$C=\frac{1}{Q}\sum_{q=1}^{Q}\mathbf{1}\Bigl{(}\hat{y}_{1,q}=\hat{y}_{2,q}=\cdots=\hat{y}_{L,q}\Bigr{)}$
4.3 Experimental Results
4.3.1 AdaCoT improves multilingual factual reasoning
Model en zh id
LLaMA3.1-8B 8.33 7.53 7.51
LLaMA3.1-8B-AdaCoT 8.35 8.13 8.13
Qwen2.5-7B 8.58 8.47 7.85
Qwen2.5-7B-AdaCoT 8.58 8.53 8.16
Table 2: CrossAlpaca-Eval 2.0: Multilingual Performance Comparison of LLaMA 3.1-8B and Qwen2.5-7B with and without AdaCoT.
CrossMMLU CrossLogiQA
Model en zh id Cons. en zh id Cons.
LLaMA3.1-8B-Inst 82.0 68.0 67.3 66.7 61.4 54.5 46.6 51.1
LLaMA3.1-8B-AdaCoT 82.6 74.6 68.0 70.6 64.8 56.3 47.7 55.7
Qwen2.5-7B-Inst 84.0 80.7 74.7 73.3 69.3 75.0 55.1 59.1
Qwen2.5-7B-AdaCoT 84.0 83.3 78.0 76.0 73.9 74.4 59.7 70.5
Table 3: Multilingual performance comparison on CrossMMLU and CrossLogiQA. Consistency is based on the predicted answers across three languages.
Table 1 reports the comparison results on multilingual TruthfulQA dataset. The LLaMA3.1-8B-AdaCoT model demonstrates performance improvements across all primary languages. The model achieves a 2.5%, 5.7% and 7.2% relative performance gain in English, Chinese and Indonesian respectively compared to the base LLaMA3.1-8B model. We also find the AdaCoT enhanced LLaMA model generalises to other low-resource languages, including Danish, Basque, Armenian, and Russian, with performance improvements exceeding 9% in absolute terms. The LLaMA3.1-8B-AdaCoT model shows marked improvements in 30 out of 32 languages tested. For Qwen 2 7B, AdaCoT yields more modest improvements in primary languages with a relative performance increase that is smaller compared to the LLaMA model, possibly due to the much smaller performance gaps across all primary languages.
Table 2 presents the evaluation results on the CrossAlpaca-Eval 2.0 dataset. We first observe that both LLaMA3.1-8B and Qwen2.5-7B base models perform strongly in English, with Qwen2.5-7B also demonstrating strong performance in Chinese. Then we find the AdaCoT method maintains the same performance in English while exhibiting significant performance gains in Chinese and Indonesian for the LLaMA3.1-8B model. It also yields small performance gains in Chinese and noticeable performance gains in Indonesian for Qwen2.5-7B model. The results show the effectiveness of AdaCoT which leverages the knowledge from high-resource languages to improve response quality of general instructions in low-resource languages.
We also notice similar performance gains in reasoning benchmarks like CrossMMLU and CrossLogiQA dataset as shown in table 3. For CrossMMLU dataset, the AdaCoT method improves both Qwen2.5-7B and LLaMA3.1-8B in Chinese and Indonesian reasoning while maintaining same performance in English. For CrossLogiQA dataset, we find the AdaCoT method also improves English which is already a high resource language, suggesting the adaptive language routing can also benefit high resource language settings. Finally, we find that AdaCoT improves the consistency of the models’ responses when they are prompted with the same question in different primary languages.
CrossAlpaca-Eval 2.0
Model en zh id
LLaMA3.1-8B (Baseline) 8.33 7.53 7.51
LLaMA3.1-8B-AdaCoT (Direct) 8.35 7.38 7.59
LLaMA3.1-8B-AdaCoT (English) 8.35 7.71 7.86
LLaMA3.1-8B-AdaCoT (Chinese) 6.99 7.38 7.23
LLaMA3.1-8B-AdaCoT (Indonesian) 7.11 7.21 7.59
LLaMA3.1-8B-AdaCoT (w/o Filter) 8.23 8.07 8.11
LLaMA3.1-8B-AdaCoT 8.35 8.13 8.13
Table 4: Ablation study results on CrossAlpaca-Eval 2.0. We report the mean GPT-4o scores for English (en), Chinese (zh), and Indonesian (id). “Direct” indicates no intermediate reasoning, while “English,” “Chinese,” or “Indonesian” indicates intermediate reasoning in that language. “w/o Filter” applies adaptive routing but omits score-based filtering.
4.3.2 Adaptive Language Routing for Enhanced AdaCoT Performance
Adaptive Language Routing (ALR) serves as a fundamental mechanism in the AdaCoT framework, which improves multilingual factual reasoning by strategically selecting intermediate thinking language based on characteristics of input prompt and performance feedback. ALR enables a contextually adaptive reasoning process that can draw upon the most suitable pre-trained linguistic knowledge for each unique instruction.
Table 5: Case Studies on Adaptive Multilingual Reasoning in the AdaCoT Framework
Input Query Pathway LLaMA3.1-8B-AdaCoT LLaMA3.1-8B
创作一首秋天的诗歌 (Compose a poem about autumn) Direct Generation 秋天的风,带来凉爽的气息; 落叶的黄金,铺满了大地的面貌 秋天的叶子飘落下,金黄的色彩在空中舞。风轻轻地吹过,带来凉爽,秋天的美景,永远留在心头。
I’m a word that rhymes with ’toast’ and means to brag or show off. What am I? Direct Generation You’re "boast"! The word that rhymes with ’toast’ and means to brag or show off is ’boast’.
抛两次硬币,至少有一个正面的概率是多少? (Flip a coin twice. What is the probability of getting at least one heads?) Intermediate Reasoning in English 第一步:确定抛两枚硬币时的总结果数量。 第二步:列出抛两枚硬币的所有可能结果。 第三步:识别至少有一枚硬币落在正面的结果。 第四步:计算至少有一枚硬币落在正面的概率。 有 3 种结果,其中至少有一枚硬币落在正面,总共有 4 种可能的结果。因此,概率为 3/4。 最终答案是:3/4。 抛两次硬币,总共有 4 种可能的结果:HH、HT、TH、TT
其中 H 代表正面,T 代表反面。至少有一个正面的概率是 1 - 至少一个反面的概率。至少一个反面的概率是:P(HT) + P(TH) + P(TT) = 1/4 + 1/4 + 1/4 = 3/4。因此,至少有一个正面的概率是:1 - 3/4 = 1/4。所以,抛两次硬币,至少有一个正面的概率是 1/4。
Jalan tol mana di Singapura yang terpanjang? (Which expressway in Singapore is longest?) Intermediate Reasoning in English Jalan tol terpanjang di Singapura adalah Pan-Island Expressway (PIE). Panjangnya sekitar 44 kilometer (27 mil) dan menghubungkan bagian barat dan timur pulau tersebut. Jalan tol terpanjang di Singapura adalah Kallang-Paya Lebar Expressway (KPE).
We conduct an ablation study on CrossAlpaca-Eval 2.0 to systematically analyze the influence of different language routing strategies. This dataset allows us to compare performance of the models across more input instruction types and more comprehensively explore the various routing strategies. The ablation study compares five variants of our approach applied to the LLaMA 3.1-8B model, alongside the official LLaMA 3.1-8B baseline and the original AdaCoT model. Table 4 summarizes the mean GPT-4o scores in all three primary languages.
We first investigate the effect of consistently “thinking” in a single auxiliary language by analyzing four variants against the official LLaMA 3.1-8B baseline and then analyses the effectiveness of score-based filtering method. We observe that: (1) Thinking exclusively in English consistently outperforms thinking in other primary languages, which is largely attributable to the model’s extensive English pre-training. This finding underscores English as the dominant knowledge resource for diverse tasks within the LLaMA model. (2) The score-based filtering mechanism can improve the performance of ALR, which demonstrates the importance of retaining only high-quality training examples. (3) Limiting thinking to English is significantly worse than adaptive language routing. The results show 5.2% and 9.3% relative performance gains compared to thinking only in English, suggesting the need to dynamically adjust reasoning strategies for improved performance.
4.3.3 Case Studies on AdaCoT
Table 5 shows how LLaMA3.1-8B-AdaCoT adapts its reasoning pathways based on the prompt. In linguistically dependent tasks like composing a poem, the model strategically generates content directly in Chinese, leveraging the language’s inherent semantic richness to preserve poetic fluency and avoid the potential information loss associated with translation or intermediate reasoning processes. Likewise, when prompted with an English word riddle that simply asks for a rhyming word, AdaCoT again employs direct generation and provides the correct answer on par with the baseline. Notably, in both cases, the AdaCoT fine-tuned model shows no degradation in answer quality.
For questions where intermediate reasoning can help, AdaCoT leverages high-resource languages (e.g., English) to boost accuracy. An example is the Chinese probability question, where the base LLaMA3.1-8B model incorrectly predicts the chance of getting at least one head in two coin tosses, but AdaCoT’s chain-of-thought in English yields the correct $3/4$ answer in Chinese. Similarly, when asked in Indonesian about Singapore’s longest expressway, the baseline mistakenly identifies the KPE, while AdaCoT correctly names the PIE by tapping into its richer English-based knowledge. These examples underscore the AdaCoT’s adaptive reasoning approach, highlighting its capacity to dynamically select optimal linguistic pathways and significantly improve multilingual reasoning across diverse task domains while maintaining performance in high-resource language setting.
5 Conclusion
We introduced AdaCoT, a framework that strategically uses intermediate “thinking languages” to boost cross-lingual factual reasoning. By adaptively routing queries through well-resourced or better-aligned languages, AdaCoT bridges performance gaps in low-resource languages while retaining strong results in high-resource settings. We also introduce an AdaCoT fine-tuning process that leverages a reward LLM to rank different reasoning pathways and selects the best pathway to optimize the model. Our experiments on Multilingual TruthfulQA, CrossAlpaca-Eval 2.0, CrossMMLU, and CrossLogiQA show consistent improvements, highlighting the benefits of Adaptive Language Routing in selecting the optimal reasoning pathway.
Limitation
While our experiments show the potential of leveraging three primary thinking languages to enhance cross-lingual factual reasoning, the current approach has limitations in comprehensive linguistic transfer. The framework depends on a limited set of languages which restricts its generalization across broader linguistic contexts. Expanding the number of thinking languages presents a complex trade-off: potential performance improvements in underrepresented languages are counterbalanced by substantive challenges, including increased training complexity and the risk of introducing semantic distortions from wrongly translated data. Furthermore, the adaptive routing mechanism, despite its innovative design, introduces computational inefficiencies manifested through higher inference latency. Lastly, AdaCoT requires diverse, high-quality training instructions which may be hard to obtain for certain domains or low-resource languages.
Acknowledgment
This research is supported by (a) National Research Foundation, Singapore and Infocomm Media Development Authority, Singapore under its National Large Language Models Funding Initiative and (b) the resources and platforms provided by the Google’s TPU Research Cloud (TRC) and Singapore NSCC Aspire2A+.